OpenStudy (anonymous):

If : \(\color{blue}{\text{2 + 6 + 10 + 14 + 18 + 22 + 26 + 30 + 34 + 38 = 200}}\), Then, \(\color{red}{\text{Add}}\) any \(\color{cyan}{\text{Five}}\) numbers and get \(\color{green}{100}\)..

OpenStudy (anonymous):

38-34x26+6-10=100 expl: 38-34=(4)x26=(104)+6=(110)-10=(100)